# GDI-Bench: A Benchmark for General Document Intelligence with Vision and Reasoning Decoupling

### v1

@@ 3.4 Comparison with Existing Document Intelligence Benchmarks
As shown in Table 2, most existing document intelligence benchmarks are limited in either scenario coverage, task diversity, or data scale. In contrast, the GDI-Bench offers a balanced combination of 9 diverse document scenarios and 19 representative tasks, based on a curated set of 1.9k images. Notably, it introduces difficulty grading , a feature absent in all other compared benchmarks.
@@
Full text extraction of sliced newspaper pages.
Extract header information of newspaper pages.……
@@
Extract full-text page content and organize it in markdown format.
Answer reasoning questions for a paragraph on a page scientific paper page (/ multiple choice).……
@@
Extract the content of the exam paper based on the type of question and the question number.
@@
4 Methodology
(a) The mean values of parameter changes per layer before and after full-parameter SFT.
(b) Distribution of parameter differences before and after full-parameter SFT.
GDI-Bench presents unique challenges for model designs, motivating the application of supervised fine-tuning to improve task-specific performance. To tackle this, we develope a multi-source training set that includes tasks structurally aligned with the benchmark, but drawn from diverse data domains, as shown in Table 3. We perform SFT on the InternVL3-8B model using this training set. However, subsequent evaluation of the model reveals significant catastrophic forgetting, wherein the model loses essential knowledge acquired during pre-training.
To investigate the phenomenon, we conduct a parameter update analysis of models undergoing full-parameter supervised fine-tuning. Our experimental results reveal significant sparsity in learned parameter adaptations: over 95% of parameters exhibit minimal changes, while a critical sparse subset (5%) undergoes substantial modification (>0.005), as depicted in Fig. 4(b). This finding aligns with and extends the Lottery Ticket Hypothesis [51], leading to our formal hypothesis: For any pre-trained MLLM with parameters $\Theta=\{\theta_{i}\}_{i=1}^{n}$, there exists a sparse task-salient sub-network $\hat{\Theta}=\{\hat{\theta}_{j}\}_{j=1}^{m}$ ($m<n$) and a domain-specific transformation $\psi$, such that $\psi(f_{\hat{\Theta}}(x))$ achieves performance comparable to the full model $f_{\Theta}(x)$ on the target domain $\mathcal{D}$.
where $\mathcal{P}(\cdot)$ is the metric for evaluating the performance of the model, and the subnetwork $\hat{\Theta}$ consists of parameters exhibiting high update sensitivity during SFT. Since only a small subset of parameters is relevant to task-specific adaptation, we propose performing gradient updates solely on these parameters during SFT, while freezing the majority of the remaining parameters to preserve the model’s generalization and essential knowledge as much as possible.
We formalize this approach as Layer-wise Adaptive Freeze-Tuning (LW-AFT), which identifies and updates only critical parameters through layer-wise sensitivity analysis. By maintaining stable base network dynamics and avoiding large-scale parameter shifts, LW-AFT mitigates catastrophic forgetting while enabling efficient adaptation to downstream tasks. The overview of our method is shown in Fig. 5. For a specific new domain $\mathcal{D}_{t}$, the domain-specific expert model (denoted as $\mathcal{E}_{t}$) is fine-tuned based on the pre-trained parameters $\Theta$, using only a fraction $\frac{1}{\alpha}$ of the full dataset. For each domain-specific expert model, the element-wise absolute change in parameters is computed as:
We perform a comprehensive layer-wise analysis of parameter updates $\Delta\theta^{(t)}_{i}$ before and after SFT. Fig. 4(a) illustrates the average magnitude of these updates per layer, revealing distinct trends between the vision and language components of the model. Specifically, vision layers exhibit more parameter modifications compared to their language layers. Intuitively, we assume that the layers undergoing substantial parameter updates are critical for domain adaptation, as their dynamics closely align with task-specific knowledge transfer. Conversely, layers with minimal updates appear less specialized to domain-specific features, and preserving their stability supports cross-domain generalization.
Figure 5: Overview of the Layer-wise Adaptive Freeze-Tuning method.
Therefore, we implement a layer-adaptive unfrozen parameter allocation strategy. Partitioning the model into $L$ architectural layers with parameter counts $W=\{w_{1},...,w_{L}\}\subseteq\mathbb{R}$, the distribution of trainable parameters across layers follows a proportional allocation mechanism: Given a global budget of $H$ unfrozen parameters, each layer $\ell$ receives a number of unfrozen parameters $h_{\ell}$ determined by
$h_{\ell}=\frac{m_{\ell}\cdot w_{\ell}}{\sum_{i=1}^{L}m_{i}\cdot w_{i}}\cdot H$ (3)
After determining the number of unfreezing parameters for each layer $\ell$, we first sort the absolute changes $\{\Delta\theta_{j}^{(t)}\}_{\ell}$ in descending order, and then select only the top $h_{\ell}$ parameters for further updates. This selection is performed via the function:
@@ 4 Methodology
The binary masks obtained for the expert $\mathcal{E}_{t}$ is denoted as $\mathcal{M}_{\mathcal{E}_{t}}=\left\{\phi_{h_{\ell}}\bigl{(}\{\Delta\theta_{j}^{(t)}\}_{\ell}\bigr{)}\right\}_{\ell=1}^{L}$. In the subsequent fine-tuning on a target domain, the gradient update is modified so that only the parameters with $\mathcal{M}_{\mathcal{E}_{t}}$ are updated:
where $\odot$ denotes the Hadamard product. This approach not only reduces the number of parameters to be trained but also preserves full-domain generalization without resorting to experience replay or other sophisticated fine-tuning techniques.
5 Experiments
The experiments in this section consist of two parts. In Section 5.1, we discuss the impact of our proposed Layer-wise Adaptive Freeze-Tuning (LW-AFT) method on cross-domain and cross-task generalization. In Section 5.2, we perform a comprehensive benchmark evaluation on the GDI-Bench.
Model T1 T2 T3 T4
date editor email phone
InternVL 3-8B 0.216 0.383 0.840 0.101 0.884 0.398 0.398
LoRA Fine-Tuning 0.102 0.473 0.300 0.093 0.606 0.171 0.172
LW-AFT (Ours) 0.096 0.365 0.200 0.010 0.316 0.058 0.153
5.1 Layer-wise Adaptive Freeze-Tuning.
5.1.1 Cross-domain and cross-task evaluation.
To validate the generalization capability of the LW-AFT method in cross-domain and cross-task scenarios, we designed the following experiments.
Cross-domain experiments with the same task.
  • T1: Train the paragraph starting position OCR task in the academic paper domain and test the same task in the Infograph domain.
  • T2: Train the information formatting and organization (JSON) task in the exam and academic paper domains, and test it in the Infograph domain.
  • T3: Train the reasoning and question-answering task in the academic paper domains, and test it in the financial report domain.
Cross-task experiments with the same domain.
  • T4: In the newspaper domain, only train the header information extraction and its formatting organization (JSON) task, and test the extraction task for header detail information (including extraction of newspaper date, editor, email, and phone number).
We compare the performance of full-parameter fine-tuning, LoRA fine-tuning, and our method under both settings, as shown in Table 4, which displays the edit distance for each task. As can be seen, our method demonstrates strong cross-domain and cross-task capabilities, significantly outperforming the LoRA fine-tuning.
5.1.2 Performance on general datasets.
We evaluate the performance of full-parameter fine-tuning, LoRA fine-tuning, and our method on a range of general datasets, as shown in Table 5. It can be seen that our method preserved the model’s general capabilities and avoided catastrophic forgetting after SFT.
Table 5: The performance of different training methods on general datasets.
[…]
val test test no mask test - val - en cn
InternVL 3 8B 92.0 86.6 85.2 92.6 80.2 880 75.6 69.7 85.5 85.6
Full-Parameter Fine-Tuning 39.4 33.3 43.1 49.4 22.7 228 25.9 44.8 15.6 14.6
LoRA Fine-Tuning 91.3 85.8 85.4 93.1 81.6 862 74.4 69.7 85.2 84.8
LW-AFT (Ours) 91.6 85.8 85.6 93.2 82.1 871 75.8 70.6 85.2 84.8
Table 6: The performance of different training methods on general datasets.
Model R0V0 R0V1 R0V2 R1V0 R1V1 R1V2 R2V0 R2V1 R2V2 Overall
Claude3-5V-Sonnet 0.10 0.74 0.78 0.60 0.52 0.91 0.18 0.52 0.62 0.58
Gemini-2.0-Flash 0.05 0.42 0.52 0.62 0.25 0.55 0.05 0.31 0.69 0.39
GPT4o-2024-11-20 0.04 0.40 0.57 0.59 0.40 0.58 0.12 0.48 0.68 0.43
Grok-2-Vision 0.33 0.80 0.84 0.80 0.71 0.70 0.23 0.54 0.62 0.66
Qwen2.5-VL-72B-Instruct 0.03 0.35 0.44 0.41 0.24 0.33 0.11 0.36 0.50 0.31
GDI-Model 0.03 0.35 0.44 0.14 0.21 0.24 0.08 0.23 0.34 0.25
5.2 Benchmark Evaluation Results.
To assess the effectiveness of our GDI model on generalized OCR tasks, we compare it against several state-of-the-art vision-language models—namely Qwen2.5-VL-72B, Gemini-2.0-Flash, GPT-4o-2024-11-20, Claude-3-5-Sonnet, and Grok-2-Vision—on GDI-Bench’s multi-level challenge suite (v0-r0 through v2-r2). All evaluations were conducted with the same preprocessing pipeline.
On v0-r0, our 8B model matches or slightly exceeds performance of Qwen2.5-VL-72B. On the held-out generalization set, GDI-Model retains a slight advantage over the much larger models (70 B+), demonstrating superior resource-efficiency and robustness. These results confirm that, despite its smaller parameter count, our model achieves performance on par with or exceeding that of much larger open- and closed-source competitors.
6 Conclusion
We introduce the GDI-Bench, a document-domain benchmark with extensive domain coverage that pioneers a systematic difficulty grading system. Utilizing a complexity decoupling mechanism, the GDI-Bench breaks down multimodal document tasks into two orthogonal dimensions: visual complexity and reasoning complexity. It establishes a quantifiable, hierarchical standard that aids in identifying model issues and guiding optimization efforts. In addition, we propose the Layer-wise Adaptive Freeze-Tuning method and the corresponding GDI model. LW-AFT mitigates the problem of catastrophic forgetting during SFT, retains 99% of the model’s general capabilities during the SFT process, and enhances its performance on cross-domain and cross-task scenarios. The GDI model has demonstrated outstanding performance across multiple document-domain benchmarks, including the GDI-Bench, as well as general datasets. We hope that the GDI-Bench can help the advancement of future models and perhaps inspire new theoretical insights.

### v2

@@ 3.4 Comparison with Existing Document Intelligence Benchmarks
As shown in Table 2, most existing document intelligence benchmarks are limited in either scenario coverage, task diversity, or data scale. In contrast, the GDI-Bench offers a balanced combination of 9 diverse document scenarios and 19 representative tasks, based on a curated set of 2.3k images. Notably, it introduces difficulty grading, a feature absent in all other compared benchmarks.
4 Methodology
Figure 5: Distribution of parameter differences before and after full-parameter SFT.
Figure 6: The mean values of parameter changes per layer before and after full-parameter SFT using 10% of the training set.
We evaluated the InternVL3-8B model on GDI-Bench, conducting a detailed weakness analysis. As shown in Fig. 2, the model performs well on R0 but drops significantly on R1 and R2, highlighting a large performance gap. To validate the ability of GDI-Bench to identify model weaknesses, we aimed to address these gaps in R1 and R2. For this purpose, we constructed a multi-source training dataset that is structurally aligned with the tasks in R1 and R2, but sourced from entirely different data domains, detailed in App. A.2. We performed SFT on the InternVL3-8B model using the constructed training set. However, subsequent evaluations revealed significant catastrophic forgetting, where the model lost important knowledge acquired during pretraining. As shown in Table 5, the Full-Parameter Fine-Tuning Model exhibited degradation in the R0 domain, where it had previously performed well.
To investigate the phenomenon, we comprehensively analyze parameter update dynamics in models subjected to full-parameter supervised fine-tuning. Our experimental results reveal significant sparsity in learned parameter adaptations: over 95% of parameters exhibit minimal changes, while a critical sparse subset (5%) undergoes substantial modification (>0.005), as depicted in Fig. 6. Inspired by the Lottery Ticket Hypothesis [52], we integrate this discovery with the hypothesis and formulate our theoretical proposition: For any pre-trained MLLM with parameters $\Theta=\{\theta_{i}\}_{i=1}^{n}$, there exists a sparse task-salient sub-network $\hat{\Theta}=\{\hat{\theta}_{j}\}_{j=1}^{m}$ ($m<n$) and a domain-specific transformation $\psi$, such that $\psi(f_{\hat{\Theta}}(x))$ achieves performance comparable to the full model $f_{\Theta}(x)$ on the target domain $\mathcal{D}$.
where $\mathcal{P}(\cdot)$ is the metric for evaluating the performance of the model, and the subnetwork $\hat{\Theta}$ consists of parameters exhibiting high update sensitivity during SFT. Given that only a small subset of parameters exhibit significant task-specific adaptation, we propose performing gradient updates solely on these parameters during SFT, while freezing the majority of remaining parameters to maximally preserve the model’s generalization capabilities and foundational knowledge.
We formalize this approach as Layer-wise Adaptive Freeze-Tuning (LW-AFT), which identifies and updates only critical parameters through layer-wise sensitivity analysis. The overview of our method is shown in Fig. 7. For a specific new domain $\mathcal{D}_{t}$, the expert model (denoted as $\mathcal{E}_{t}$) is fine-tuned based on the pre-trained parameters $\Theta$ using only a subset $\alpha$ ($\alpha<1$) of the full dataset. For each domain-specific expert model, the element-wise absolute parameter variation is computed as:
We conduct a comprehensive layer-wise analysis of parameter updates $\Delta\theta^{(t)}_{i}$ before and after SFT. Fig. 6 illustrates the average magnitude of updates per layer, revealing distinct trends between the vision and language components. Specifically, the language layers exhibit a higher degree of parameter modification compared to the visual layers. This observation aligns with the problem localization we performed using GDI-Bench, where the InternVL3-8B model shows a performance degradation in reasoning dimensions. Intuitively, we assume that the layers undergoing substantial parameter updates are critical for domain adaptation, as their dynamics closely align with task-specific knowledge transfer. Conversely, layers with minimal updates appear less specialized to domain-specific features, and preserving their stability supports cross-domain generalization.
Figure 7: Overview of the Layer-wise Adaptive Freeze-Tuning method.
Therefore, we implement a layer-adaptive freezing parameter allocation strategy. The model is divided into $L$ architectural layers with parameter counts $W=\{w_{1},...,w_{L}\}\subseteq\mathbb{R}$. Given a global budget of $H$ unfrozen parameters, we employ a proportional allocation mechanism where each layer $\ell$ is allocated $h_{\ell}$ unfrozen parameters, computed as follows:
$h_{\ell}=\frac{m_{\ell}\cdot w_{\ell}}{\sum_{i=1}^{L}m_{i}\cdot w_{i}}\cdot H$ (5)
After determining the number of unfrozen parameters for each layer $\ell$, we first sort the absolute changes $\{\Delta\theta_{j}^{(t)}\}_{\ell}$ in descending order, and then select only the top $h_{\ell}$ parameters for further updates. This selection is performed via the function:
[…]
The binary masks for the entire network of expert $\mathcal{E}_{t}$ can be formulated as $\mathcal{M}_{\mathcal{E}_{t}}=\left\{\phi_{h_{\ell}}\bigl{(}\{\Delta\theta_{j}^{(t)}\}_{\ell}\bigr{)}\right\}_{\ell=1}^{L}$. During the subsequent fine-tuning on a target domain, gradient updates are dynamically masked to exclusively modify parameters corresponding to $\mathcal{M}_{\mathcal{E}_{t}}$ through:
where $\odot$ denotes the Hadamard product. By maintaining the stability of the base network dynamics and avoiding large parameter shifts, LW-AFT effectively mitigates catastrophic forgetting while adapting efficiently to downstream tasks. This approach not only reduces the number of parameters that need to be trained but also preserves global generalization capabilities, without relying on experience replay or other complex fine-tuning techniques. With the LW-AFT method, we are able to retain the outstanding performance of the InternVL3-8B model in the R0 domain, while improving its performance in the R1 and R2 domains, as shown in Table 5. This further demonstrates the ability of GDI-Bench to pinpoint model weaknesses.
5 Experiments
The experiments in this section consist of two parts. In Section 5.1, we discuss the impact of our proposed Layer-wise Adaptive Freeze-Tuning (LW-AFT) method. In Section 5.2, we perform a comprehensive benchmark evaluation on the GDI-Bench. All experiments are based on the InternVL3-8B model [53], with detailed training parameters provided in App. A.1.
5.1 Layer-wise Adaptive Freeze-Tuning.
5.1.1 Ablation Study.
Figure 8: Ablation study on the $\alpha$.
Figure 9: Ablation study on the freeze rate of parameters.
We perform an ablation study on the parameter $\alpha$ in Section 4, which controls the fraction of the dataset used to train domain-specific expert models. Fig. 9 presents the results of this ablation study, illustrating how expert models trained with different $\alpha$ values influence the final performance on GDI-Bench. In all settings, the LW-AFT method is used with 99% of the model parameters frozen. The results indicate that $\alpha=10\%$ achieves the best performance and data efficiency, and is therefore adopted in our default configuration.
We conduct an ablation study to assess the impact of varying the proportion of frozen parameters in the model. Fig. 9 illustrates the performance of the 8B-sized model across GDI-Bench and several other datasets under different freeze ratios. After evaluating the trade-off between maintaining strong performance on GDI-Bench and ensuring generalization across the additional datasets, we determine that freezing 99% of the model parameters represents the optimal configuration for our final setup.
5.1.2 Evaluation of Catastrophic Forgetting Severity.
To validate the effectiveness of LW-AFT in mitigating catastrophic forgetting, we evaluate the performance of Full-Parameter Fine-Tuning, LoRA Fine-Tuning, and the LW-AFT method across a range of datasets, including OCR, Chart, Document Understanding, and Multimodal Multilingual Understanding datasets [54, 16, 55, 56, 57, 19, 58, 10, 59], as shown in Table 3. The results demonstrate that our method prevents catastrophic forgetting after SFT. Furthermore, on certain datasets, such as TextVQA, InfoVQA, and Seed2plus, our method outperforms the base model, which validates the efficacy of our intelligence-preserving training strategy.
Table 3: The performance of different training methods on multiple datasets.
[…]
val test test no mask test - - val - en cn
Base 92.0 86.6 85.2 92.6 80.2 880 0.21 75.6 69.7 85.5 85.6
Full-Parameter Fine-Tuning 39.4 33.3 43.1 49.4 22.7 228 0.42 25.9 44.8 15.6 14.6
LoRA Fine-Tuning 91.3 85.8 85.4 93.1 81.6 862 0.23 74.4 69.7 85.2 84.8
LW-AFT (Ours) 91.6 85.8 85.6 93.2 82.1 871 0.23 75.8 70.6 85.2 84.8
5.1.3 Cross-domain and Cross-task Evaluation.
To validate the generalization capability of the LW-AFT method in cross-domain and cross-task scenarios, we designed the experiments covering the following two types of tasks.
Cross-domain experiments with the same task.
  • T1: The model is trained for the paragraph beginning extracting task in the Scientific paper domain, and evaluated on the same task in the Infograph domain.
  • T2: The model is trained for the information extraction and format organization task in the Exam paper domains, and evaluated on the same task in the Infograph domain.
  • T3: The model is trained for the reasoning and question-answering task for tables in the scientific paper domain, and evaluated on the same task in the Financial report domain.
Cross-task experiments with the same domain.
  • T4: In the Newspaper domain, the model is trained for header information extraction and format organization tasks, and evaluated on fine-grained header field extraction performance (including publication date, editor information, email address, and phone number extraction).
Method T1 T2 T3 T4
date editor email phone
Base 0.216 0.383 0.840 0.101 0.884 0.398 0.398
LoRA Fine-Tuning 0.102 0.473 0.300 0.093 0.606 0.171 0.172
LW-AFT (Ours) 0.096 0.365 0.200 0.010 0.316 0.058 0.153
We compare the performance of Full-Parameter Fine-Tuning, LoRA Fine-Tuning, and the LW-AFT method under both settings, as shown in Table 4, which displays the normalized edit distance for each task. As can be seen, our method demonstrates strong cross-domain and cross-task capabilities, significantly outperforming the LoRA fine-tuning, further demonstrating the effectiveness of the LW-AFT method’s generalization ability in the task dimension.
Table 5: The performance of different open-source and closed-source large models and different training methods on GDI-Bench.
Model Size R0V0 R0V1 R0V2 R1V0 R1V1 R1V2 R2V0 R2V1 R2V2 Overall
Claude-3-5-Sonnet (2024-10-22) 175B 0.90 0.26 0.22 0.42 0.63 0.20 0.86 0.59 0.44 0.509
Gemini-2.0-Flash 750B 0.94 0.58 0.48 0.37 0.77 0.69 0.92 0.73 0.41 0.662
GPT-4o (2024-11-20) 200B 0.96 0.60 0.43 0.39 0.71 0.57 0.81 0.51 0.35 0.593
Grok-2-Vision-on 500B 0.67 0.20 0.16 0.21 0.31 0.30 0.77 0.46 0.41 0.377
Qwen2.5-VL-72B-Instruct 72B 0.97 0.69 0.57 0.60 0.81 0.72 0.90 0.63 0.49 0.706
InternVL3-8B 8B 0.96 0.65 0.56 0.35 0.77 0.47 0.89 0.62 0.50 0.652
Full-Parameter Fine-Tuning Model 8B 0.94 0.56 0.39 0.64 0.71 0.67 0.65 0.44 0.51 0.599
LoRA Fine-Tuning Model 8B 0.97 0.66 0.56 0.87 0.82 0.77 0.91 0.71 0.68 0.759
GDI-Model (Ours) 8B 0.97 0.65 0.56 0.89 0.82 0.77 0.93 0.74 0.68 0.762
5.2 Benchmark Evaluation Results.
We employ the LW-AFT method on the complete training dataset, resulting in the development of an 8B-sized GDI-Model. To assess the effectiveness of our GDI-Model on generalized OCR tasks, we compare it against several SOTA vision-language models—namely Qwen2.5VL-72B [3], Gemini-2.0-Flash [5], GPT-4o (2024-11-20) [60], Claude-3-5-Sonnet, and Grok-2-Vision-on and InternVL3-8B [53] on GDI-Bench’s multi-level challenge suite (V0-R0 through V2-R2). All evaluations were conducted with the same preprocessing pipeline, as shown in Fig. 2 and Table 5.
Compared to the base InternVL3-8B, our model maintains strong R0-level performance and shows clear gains at R1 and R2, validating the effectiveness of the LW-AFT method. GDI-Model matches Qwen2.5-VL-72B at R0 and outperforms it at higher reasoning levels. At R2, our model rivals Gemini-2.0-Flash under low to medium visual complexity (V0/V1), and even surpasses it under high complexity (V2). These results highlight that, despite a smaller parameter size, our model achieves or exceeds the performance of much larger open- and closed-source counterparts.
6 Conclusion
We introduce GDI-Bench, a document-domain benchmark with broad coverage and a pioneering difficulty grading system. By decoupling visual and reasoning complexity, it enables systematic evaluation and model optimization guidance. Alongside, we introduce LW-AFT (Layer-wise Adaptive Freeze-Tuning) and the corresponding GDI-Model, which reduces catastrophic forgetting during SFT. This method preserves general capabilities while boosting cross-domain and cross-task performance. GDI-Model achieves strong results on GDI-Bench and other benchmarks. We hope that the GDI-Bench can help the advancement of future models and perhaps inspire new theoretical insights.
Appendix A Technical Appendices and Supplementary Material
A.1 Training Details and Hyperparameters
We use the InternVL3-8B model as the base model and perform supervised fine-tuning on it. In our experiments, both the LW-AFT method and Full-Parameter Fine-Tuning utilize the same set of training hyperparameters, as shown in Table 6. The training process was conducted on 8 A100 GPUs. The specific hyperparameters for LoRA fine-tuning are provided in Table 7.
Table 6: Hyperparameters of the Full-Parameter Fine-Tuning and Layer-wise Adaptive Freeze-Tuning.
Hyperparameters Value
Force image size 448
Drop path rate 0.1
Max dynamic patch 20
Train epochs 1
Learning rate 3e-5
Weight decay 0.05
Warm-up ratio 0.03
Lr scheduler cosine
Batch size 12
Max seq length 8192
Table 7: Hyperparameters of the LoRA Fine-Tuning.
Hyperparameters Value
Force image size 448
Drop path rate 0.0
Use LLM LoRA 16
Max dynamic patch 20
Train epochs 1
Learning rate 3e-5
Weight decay 0.05
Warm-up ratio 0.03
Lr scheduler cosine
Batch size 12
Max seq length 8192
A.2 Training Data
We construct a training dataset for supervised fine-tuning to enhance the base model InternVL3-8B’s capabilities in the R1 and R2 domains. This dataset contains approximately 200,000 QA-form tasks, primarily focused on tasks similar to R1 and R2, but not identical. The data sources are strictly different from those of GDI-Bench. Detailed information on the dataset size and task settings can be found in Table 8.
@@ A.2 Training Data
Extract header information of a newspaper page.
Answer reasoning questions for newspaper page (/ multiple choice).……
@@
Answer reasoning questions for a paragraph on a scientific paper page (/ multiple choice).……
@@
Answer reasoning questions for a table on a financial report page (/ multiple choice).……
@@
A.3 Qualitative Analysis.
In this section, we provide a qualitative analysis, comparing several cases to demonstrate the performance of the GDI model, as shown in Figures 10-16.
Figure 10: Table extraction task, both the InternVL3-8B model and the Qwen2.5-VL-72B model made errors. The full parameter fine-tuned models failed to handle the table extraction task due to catastrophic forgetting and ended up performing OCR on the entire page’s content. In contrast, the GDI-Model successfully extracted the table and output it in LaTeX format.
Figure 11: Newspaper titles extraction task, both full parameter fine-tuned models and LoRA fine-tuned models tend to generate model hallucinations, whereas the GDI-Model, obtained using the LW-AFT method, is able to complete the task correctly.
Figure 12: Task of extracting and organizing test points into JSON format, the LoRA model fails to accurately locate the key points or produce correct outputs, whereas the GDI-Model, obtained using the LW-AFT method, is able to complete the task correctly.
Figure 13: The task of extracting and organizing problems by question number into JSON format, the GPT-4o model fails to correctly locate the questions based on their numbers and ends up fabricating non-existent questions. In contrast, the GDI-Model successfully extracts the questions and outputs them in the specified JSON format.
Figure 14: The task of extracting paper author information and organizing it into JSON format, the Claude3-5V-Sonnet model extracted incorrect affiliation information, while the GDI-Model successfully completed the task.
Figure 15: Table reasoning tasks in scientific paper page, the GPT-4o, Grok-2-Vision-on, and Qwen2.5-VL-72B models provided incorrect answers, whereas the GDI-Model answered correctly.
Figure 16: The task of converting chart data within a region into a markdown table, the GPT-4o model does not support the task, Qwen2.5-VL-72B produced a model hallucination, and Grok-2-Vision-on incorrectly converted the data and failed to extract it. In contrast, the GDI-Model completed the task.
A.4 Sample Prompts for Evaluation.
Prompt template for document extraction (v0 tasks) Extract all text from the image accurately without interpretation and output only the text without adding any extra explanations or comments.
Prompt template for document conversion (v1 & v2 tasks) You are an AI assistant specialized in converting PDF images to Markdown format. Please follow these instructions for the conversion: 1. Text Processing: Accurately recognize all text content in the PDF image without guessing or inferring. Convert the recognized text into Markdown format. Maintain the original document structure, including headings, paragraphs, lists, etc. 2. Mathematical Formula Processing: Convert all mathematical formulas to LaTeX format. Enclose inline formulas with \( \). For example: This is an inline formula $E=mc^{2}$. Enclose block formulas with \[ \] . For example: $\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}$ 3. Table Processing: Convert tables to Markdown format. 4. Figure Handling: Ignore figures content in the PDF image. Do not attempt to describe or convert images. 5. Output Format: Ensure the output Markdown document has a clear structure with appropriate line breaks between elements. For complex layouts, try to maintain the original document’s structure and format as closely as possible. Please strictly follow these guidelines to ensure accuracy and consistency in the conversion. Your task is to accurately convert the content of the PDF image into Markdown format without adding any extra explanations or comments.
Prompt template for newspaper date extraction task Extract the date information (only including the day, month, and year, if available) of the newspaper, preserving the original punctuation and formatting.
@@
Prompt template for single-choice questions Please answer the question according to the document, outputting the correct option letter without adding any extra explanations or comments.
Prompt template for chart extraction from document Extract raw chart data from specified PDF coordinates (x1, y1, x2, y2) and output as Markdown table with the first row fixed as header. Do not add any extra explanations or comments.
Prompt template for targeted extraction Please extract the text in the yellow color box in the document without adding any extra explanations or comments.
A.5 Limitations
In this section, we outline the limitations of GDI-Bench. Currently, GDI-Bench supports only single-image document understanding tasks, excluding multi-image or multi-document tasks. Future work will expand the benchmark to include these more complex tasks, enhancing its difficulty and enabling wider evaluations.
A.6 Broader impacts
GDI-Bench helps developers diagnose model failures in vision-reasoning coupled document understanding tasks, aiding in weakness identification and optimization. However, the decoupling analysis method could be exploited maliciously, potentially boosting the capabilities of MLLMs (Multi-modal Large Language Models) in harmful ways.